# Public-Key Quantum Money and Fast Real Transforms

### v3

@@ 8.1 Group Action Quntum Walks
Cayley graphs can also be constructed using group actions. Given a regular group action $(G,X,*)$ with a fixed element $x\in X$ and a set $Q=\{q_{1},q_{2},\dots,q_{k}\}\subset G$, let $\Gamma=(X,E)$ be a graphs with vertex set $X$ and edge set consisting of pairs $(x,y)\in X\times X$ such that $y=q*x$ for some $q\in Q$. The adjacency matrix of $\Gamma$ is
@@
Since the action $(G,X,*)$ is regular, the two constructions yield the same graph up to isomorphism. In the first graph, the vertex set is $G$, and the rows and columns of the adjacency matrix are indexed by the elements of $G$, whereas in the second graph, the vertex set is $X$, and the rows and columns of the adjacency matrix are indexed by the elements of $X$. The isomorphism between the two graphs is induced by the bijection
@@
Assuming the group action is a cryptographic group action, the isomorphism $\phi$ is a one-way function: given $g\in G$, it is easy to compute $g*x$, whereas given $(x,g*x)$, it is hard to compute $g$. Therefore, although these graphs are mathematically the same, computational problems based on them require fundamentally different techniques. A notable example is the implementation of quantum walks on these graphs. [^1] Implementations of continuous-time quantum walks in the context of cryptographic group actions have previously appeared in [8, 16], in the setting of supersingular isogeny graphs. These walks can be interpreted as group-action walks. From this point onward, assume $Q$ is a small set, i.e., $\lvert Q\rvert=\operatorname{poly}(\log\lvert G\rvert)$. In the first graph, we can efficiently implement the walks $e^{-iAt}$ even for values of $t$ that are exponentially large in $\log\lvert G\rvert$. This follows from the fact that
@@
The quantum Fourier transform $\mathsf{QFT}_{G}$ and its inverse can be applied in $\operatorname{poly}(\log\lvert G\rvert)$ operations. The diagonal unitary in the middle, which is a phase computation $\lvert a\rangle\mapsto e^{-i\lambda_{a}t}\lvert a\rangle$, is also efficient because $\lambda_{a}$ can be computed classically to arbitrary precision in $\operatorname{poly}(\log\lvert G\rvert)$ time.
@@
Despite this limitation, the sparsity and structure of the matrix $A$ allow us to demonstrate in the next section that for $t=\operatorname{poly}(\log\lvert G\rvert)$, the walk $e^{-iAt}$ can still be efficiently approximated to polynomial accuracy.
@@
Assume $t=\operatorname{poly}(\log\lvert G\rvert)$. We show that the walk $W=e^{-iAt}$ can be efficiently simulated using the discrete-time quantum walk technique from [9, 6]. Consider the isometry $T$ in (22), where the states $\lvert\phi_{b}\rangle$ are defined in (21). It suffices to show that $T$ and it inverse $T^{*}$, and the discrete-time walk $W=iS(2TT^{*}-\mathds{1})$ can be implemented efficiently for group actions. Note that for the Cayley graph $\Gamma=(X,E)$ of the group action $(G,X,*)$, the Hamiltonian $H$ is the adjacency matrix $A$ of $\Gamma$, which is indexed by the elements of $X$. The isometry $T$ then becomes
@@
We now show how to implement $T$ by constructing an efficient unitaries $U_{0}$ and $U_{1}$ defined $U_{b}:\lvert 0\rangle\lvert y,b\rangle\lvert 0,0\rangle\mapsto\lvert 0\rangle\lvert y,b\rangle\lvert\phi_{yb}\rangle$, $b=0,1$, where the first register is an ancilla. Since $\lvert\phi_{y1}\rangle=\lvert 0\rangle\lvert 1\rangle$, the unitary $U_{1}$ can be implemented efficiently. The unitary $U_{0}$ can be implemented as follows. Since $Q$ is a small set, we can construct an efficient unitary $V_{Q}:\mathbb{C}^{Q}\to\mathbb{C}^{Q}$ such $V_{Q}\lvert 0\rangle=\lvert Q\rvert^{-1/2}\sum_{q\in Q}\lvert q\rangle$. Applying the unitary $V_{Q}\otimes\mathds{1}$ to the state $\lvert 0\rangle\lvert y,0\rangle\lvert 0,0\rangle$ results in
@@
Next, we apply the unitary $V_{1}:\lvert q\rangle\lvert y,0\rangle\lvert 0,0\rangle\mapsto\lvert q\rangle\lvert y,0\rangle\lvert q*y,0\rangle$. Following this, we uncompute the first register to obtain $\lvert 0\rangle\lvert y,0\rangle\lvert\phi_{y0}\rangle$. Uncomputing the first register involves applying the unitary operation $V_{2}:\lvert q\rangle\lvert y,0\rangle\lvert q*y,0\rangle\mapsto\lvert 0\rangle\lvert y,0\rangle\lvert q*y,0\rangle$, which can be implemented efficiently since $Q$ is small and we can recover $q$ from the pair $(y,q*y)$ by checking against all the elements in $Q$.

### v4

@@ 8.1 Group Action Quantum Walks
Cayley graphs can also be constructed using group actions. Given a regular group action (G,X,)(G,X,*) with a fixed element $x\in X$ and a set $Q=\{q_{1},q_{2},\dots,q_{k}\}\subset G$, let $\Gamma=(X,E)$ be a graph with vertex set $X$ and edge set consisting of pairs $(x,y)\in X\times X$ such that y=qxy=q*x for some $q\in Q$. The adjacency matrix of $\Gamma$ is
@@
Since the action (G,X,)(G,X,*) is regular, the two constructions yield the same graph up to isomorphism. In the first graph, the vertex set is $G$, and the rows and columns of the adjacency matrix are indexed by the elements of $G$, whereas in the second graph, the vertex set is $X$, and the rows and columns of the adjacency matrix are indexed by the elements of $X$. The isomorphism between the two graphs is induced by the bijection
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
We now show how to implement $T$ by constructing efficient unitaries $U_{0}$ and $U_{1}$ defined as $U_{b}:\lvert 0\rangle\lvert y,b\rangle\lvert 0,0\rangle\mapsto\lvert 0\rangle\lvert y,b\rangle\lvert\phi_{yb}\rangle$, $b=0,1$, where the first register is an ancilla. Since $\lvert\phi_{y1}\rangle=\lvert 0\rangle\lvert 1\rangle$, the unitary $U_{1}$ can be implemented efficiently. The unitary $U_{0}$ can be implemented as follows. Since $Q$ is a small set, we can construct an efficient unitary $V_{Q}:\mathbb{C}^{Q}\to\mathbb{C}^{Q}$ such that $V_{Q}\lvert 0\rangle=\lvert Q\rvert^{-1/2}\sum_{q\in Q}\lvert q\rangle$. Applying the unitary $V_{Q}\otimes\mathds{1}$ to the state $\lvert 0\rangle\lvert y,0\rangle\lvert 0,0\rangle$ results in
@@
Next, we apply the unitary $V_{1}:\lvert q\rangle\lvert y,0\rangle\lvert 0,0\rangle\mapsto\lvert q\rangle\lvert y,0\rangle\lvert q*y,0\rangle$. Following this, we uncompute the first register to obtain $\lvert 0\rangle\lvert y,0\rangle\lvert\phi_{y0}\rangle$. Uncomputing the first register involves applying the unitary operation $V_{2}:\lvert q\rangle\lvert y,0\rangle\lvert q*y,0\rangle\mapsto\lvert 0\rangle\lvert y,0\rangle\lvert q*y,0\rangle$, which can be implemented efficiently since $Q$ is small and we can recover $q$ from the pair (y,qy)(y,q*y) by checking against all the elements in $Q$.